OpenStudy (anonymous):

What is the general form of the equation for the given circle? x2 + y2 - 8x - 8y + 23 = 0 x2 + y2 - 8x - 8y + 32 = 0 x2 + y2 - 4x - 4y + 23 = 0 x2 + y2 + 4x + 4y + 9 = 0

OpenStudy (anonymous):

wow this is hard. what grade u in?

OpenStudy (anonymous):

11 th

OpenStudy (anonymous):

you mean to say which one of these equation represents a circle right ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

okay to solve this problem we're going to use the standard form \[(y-y_0)^2+(x-x_0)^2=r^2\]

OpenStudy (anonymous):

ok so how do you do it?

OpenStudy (anonymous):

we're going to try to write your equations in this form

OpenStudy (anonymous):

if we develop the standard form we obtain \[y^2-2y_0+{y_0}^2+x^2-2x_0+{x_0}^2=r^2\]

OpenStudy (anonymous):

ok what does that mean im very confused srry

OpenStudy (anonymous):

can you figure out which one is closed to be like this form

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

this is part of the question

OpenStudy (anonymous):

what about this part

OpenStudy (anonymous):

I still don't understand :)